OpenStudy (anonymous):

The sum of two integers is 285, and the larger number is 17 more than 3 times the smaller number. Find the two integers.

OpenStudy (a_clan):

x+y=285...(1) x-3y=17....(2)

OpenStudy (anonymous):

x+y=285 y=3x+17 3x+17+x=285 4x=268 x=67, y=218